Find the potential drop along a copper wire 500 m long and 2 mm in diameter if the current in it is equal to 2 A. Determine the potential drop in resistors. ρ(copper)=1.68×10−8 (Ω·m)Energy of charged capacitor.



1
Expert's answer
2022-01-28T13:59:35-0500

"U=IR=\frac{4I\rho l}{\pi d^2}=5.3~V."